Grade 9Algebra

If x+y+z=0, then what is (x2/2 x2+yz) + (y2/2y2+xz) + (Z2/2z2+xy)?

dev
6 Years agoGrade 9

1 Answer

Arun
6 Years ago
x+y+z=0
 
x+y=-z…………(1) , cubing both sides
 
x^3+y^3+3.x.y(x+y)=-z^3 , now put x+y= -z
 
x^3+y^3–3x.y.z=-z^3
 
x^3+y^3+z^3 =3.x.y.z , divide both sides by x.y.z
 
x^2/yz+y^2/zx+z^2/xy= 3 , answer.